OpenStudy (faman39):

Henry grabs one shoe at a time at random without replacement from a box of 14 shoes in his closet. The box contains 6 brown shoes, 4 black shoes, and 4 sneakers. What is the probability that the first two shoes he chooses will be black? Round to the nearest percent if necessary. 7% 8% 16% 18%

OpenStudy (anonymous):

\[\frac{4}{14}\times \frac{3}{13}\]

OpenStudy (faman39):

Thank you so much @satellite73 ! :)

OpenStudy (lgbasallote):

wow..talk about spamming questions....

OpenStudy (faman39):

Lol

OpenStudy (faman39):

So it would be 7% nearest number